OpenStudy (ebydaqueen19):

HELP ME PLEASE !!! 0.5 GRAMS OF SODIUM CHLORIDE IS DISSOLVED TO MAKE .05 mL OF SOLUTION. WHAT IS MOLARITY ??

OpenStudy (abb0t):

molarity is moles per liters of solution. You are already given 0.5 grams of sodium chloride, so using dimensional analysis to get moles of sodium chloride. Then, divide by 0.05 mL of solution and multiply by 1000 to get Liters.

OpenStudy (ebydaqueen19):

THANK YOU :0

OpenStudy (ebydaqueen19):

I DIVIDED BY 1000..I THOUGHT TO GET FROM ML TO L YOU DIVIDE BY 1000??

OpenStudy (abb0t):

\(\sf \frac{0.05g \times molar~mass}{0.05~mL \times \frac{1~L}{1000~mL}}\)